# A Generalized Back-Door Criterion for Linear Regression

### v6

@@ 3.3 Post-Treatment Bias
In addition to the set $\bm{S}=\{X_{j}\}\cup\bm{Z}$ in equation (3.1), the following sets are introduced:
@@
The bias $\gamma_{ik\mid\bm{S}\setminus\{X_{k}\}}$ corresponding to each coefficient $\beta_{ik\mid\bm{S}\setminus\{X_{k}\}}$ in equation (3.1) is defined as
@@
Although bias is often defined as the difference from the total effect, this paper adopts the above definition for clarity in the quantitative characterization of post-treatment bias.
@@
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with equation (2.1). For any subset $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ that blocks every back-door path $X_{j}\dashleftarrow X_{i}$, the bias of the coefficient $\beta_{ij\mid\bm{Z}}$ in equation (3.1) is given by
@@ 3.4 Necessity for Almost All Parameter Points
Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with equation (2.1). Fix a set $\bm{Z}\subset\bm{V}$ and consider the coefficient $\beta_{ij\mid\bm{Z}}$ in equation (3.1). Then, except on a set of Lebesgue measure zero in the parameter space $\bm{\Theta}$, the following statements are equivalent:
@@ 3.5 Application to Nonlinear Data-Generating Processes
Figure 3: The causal path diagram associated with equation (3.3)
@@
Here, we restrict our discussion to the case of equation (3.3). Note that the derivative constituting the nonlinear effect can be expressed as follows.

### v7

@@ 3.3 Post-Treatment Bias
In addition to the set $\bm{S}=\{X_{j}\}\cup\bm{Z}$ in (3.1), the following sets are introduced:
@@
The bias $\gamma_{ik\mid\bm{S}\setminus\{X_{k}\}}$ corresponding to each coefficient $\beta_{ik\mid\bm{S}\setminus\{X_{k}\}}$ in (3.1) is defined as
@@
Although bias is often defined as the difference from the total effect, we adopt the above definition for clarity in the quantitative characterization of post-treatment bias.
@@
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any subset $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ that blocks every back-door path $X_{j}\dashleftarrow X_{i}$, the bias of the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) is given by
@@ 3.4 Necessity for Almost All Parameter Points
Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). Fix a set $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ and consider the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1). Then, except on a set of Lebesgue measure zero in the parameter space $\bm{\Theta}$, the following statements are equivalent:
@@ 3.5 Application to Nonlinear Data-Generating Processes
Figure 3: The causal path diagram associated with (3.3)
@@
Here, we restrict our discussion to the case of (3.3). Note that the derivative constituting the nonlinear effect can be expressed as follows.